OpenStudy (anonymous):

Determine the break even point of these two functions C(x)=12.50x+85.00 and S(x)=37.00x I know what the break even point i just do not know how to calculate it algebraically

OpenStudy (anonymous):

there i fixed it

OpenStudy (anonymous):

You're looking for the intersection point of the graphs of two linear equations: C(x)= y = 12.50x + 85.00 S(x) = y = 37.00x Substituting the 2nd expression for y in the 1st: 37.00x = 12.50x + 85.00 Subtract 12.5 from both sides: 24.5x = 85 x = 850/245 = 3.47 y = 37*850/245 = 128.37